The length of a rectangle is 8 times its width. The perimeter of the rectangle is 54 kilometers. Find the dimensions of the rectangle.
Answer by mananth(16946)

width = x
length = 8x
..
perimeter = 2(8x+x)=54
16x+2x=54
18x=54
x=3 km
length = 24 km.